OpenStudy (oreolover4876):

M is the midpoint of RS. R is at (4,3) and M is at (6,7). What are the coordinates of S?

OpenStudy (anonymous):

from 4 to 6 is two units right go another two and get get to ...

OpenStudy (ahsome):



OpenStudy (anonymous):

from 3 to 7 is 4 units up go another 4 and get to ?

OpenStudy (anonymous):

@Ahsome nice picture but i think you misread the question

OpenStudy (ahsome):

Whoops ;)

OpenStudy (oreolover4876):

I am honestly so confuzzled right now.

OpenStudy (ahsome):

Simply use the MidPoint Formula:

OpenStudy (ahsome):

\[\left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right)\]

OpenStudy (oreolover4876):

I am no longer confuzzled! thanks a ton!

OpenStudy (ahsome):

Where \((x_1,y_1)\) is the co-ordinate of one end point (R) and \((x_2,y_2)\) is the co-ordinate of another end point (S) Now, lets do this step by step. The section, \[\left(\frac{x_1+x_2}{2}\right)\] Tells you the \(x\) value of the midpoint. Since we know x value of the midpoint is 6, we can say that: \[\left(\frac{x_1+x_2}{2}\right)=6\] Now, we know that one of the x value for one end was 4 (R Point). So lets put that also \[\left(\frac{4+x_2}{2}\right)=6\] Now, \(x_2\) is meant to be the x value for the second end point (S), but since we do not know it, we will simply leave it as \(x\). \[\left(\frac{4+x}{2}\right)=6\] Now, if we find the value of \(x\), we will know the \(x\) value for S. So solve away! \[\left(\frac{4+x}{2}\right)=6\]\[\frac{4+x}{2}=6\]\[4+x=12\]\[x=8\] Therefore, the \(x\) value for S is \(8\). Do the same thing, but for \(y\) instead it, to get the \(y\) value for S. Here is the equation: \[\left(\frac{y_1+y_2}{2}\right)=7\]

OpenStudy (ahsome):

@Oreolover4876, that should be the steps. If you think I helped, please consider pressing the best response button Thank you :)